# Improving Grip Stability Using Passive Compliant Microspine Arrays for Soft Robots in Unstructured Terrain

### v1

@@ IV RESULTS
The translation results of 0ML, 1ML, and 2ML are compared on 4 surfaces. Fig. 10 shows the average displacement and standard deviation of the 3 trials per prototype.
@@
The concrete surface is level with a uniform distribution of asperities. Here, both 1ML and 2ML outperform 0ML in terms of displacement. In fact, 1ML had an average displacement of 39.63cm which is over 30 times as much as the 1.32cm that 0ML traveled. 2ML traveled 14.13cm, nearly 11 times as much as 0ML. 2ML had higher consistency across trials, having a standard deviation of 0.76 whereas 0ML’s standard deviation was 0.83.
@@
The partially uniform brick contains gaps in between bricks that can cause microspines to become stuck. This was observed to happen randomly across all trials. However, in every instance where this occurred, the microspine that was caught on 1ML or 2ML was able to wiggle free within a few seconds and continue moving. Even with these obstacles, both 1ML traveling 15cm and 2ML traveling 10.52cm had far greater average displacement than 0ML, roughly 25 times and 18 times more, respectively. This was the only surface where 0ML had the lowest standard deviation, and this is due to the microspine limbs on the other two robots randomly getting stuck when crossing over brick perimeters. Additionally, this surface resulted in the lowest overall movement of 0ML, with an average displacement of 0.59cm.
@@
The granular, compact sand was not entirely level and various pebbles, holes, insects, and small sticks were scattered around the testing area. This surface was difficult to overcome for all three robots as the compact sand was covered in a loose, granular top layer that was easy to become partially submerged in. Due to this, the average displacement is far lower on this surface. The two microspine limbs on 2ML seemed to dig itself deeper, resulting in average displacement of only 0.46cm and the lowest standard deviation. 1ML still outperformed 0ML with over 3 times increased displacement, traveling 2.53cm compared to 0.82cm.
@@
Figure 10: The average displacement data per prototype and surface combination. The repeatability is shown in error bars.
@@
Figure 11: Gait analysis. a) Every $\Delta$X and $\Delta$Y position per gait on concrete and brick. b) The average absolute $\Delta$X and $\Delta$Y displacement data per gait on concrete and brick.
@@
The forest floor surface was completely non-uniform with highly varying terrain. Specifically, the prototypes had to first overcome a large, 4” tall tree root and then move through leaves, acorns, and other tree debris. All SoRos were able to successfully traverse over the cylinder-like tree root due to the conformable nature of the soft limbs, highlighted in Fig. 2. However, only the 1ML and 2ML were able to navigate through the thick tree debris after making it over the large tree root; 0ML became stuck at the base of the tree root in each of its three trials. This is exhibited in the average displacement of 13.74cm for 0ML, 23.95cm for 1ML, and 21.32cm for 2ML. The forest floor was critical for testing as it was the most unstructured of the four experimental surfaces and showcases the benefits of the soft limbs paired with the added gripping stability of the microspine array.
@@
For each of the 36 trials, the robots move for 60 gaits, resulting in a total of 2,160 gaits or 720 gaits per robot. Each gait results in relative change in position in the robot coordinate system $\Delta X,\Delta Y$. The locomotion consistency can be analyzed by examining the 180 poses on a given surface per prototype, such as those shown in Fig. 11a). The average displacement per gait ($\Delta X,\Delta Y)$ as well as the standard deviation over all gaits per prototype/surface combination is visualized in Fig. 11b). We only analyze the uniform/partially uniform surfaces as the gait-to-gait movement is much less consistent due to non-uniformity and randomness for the other two surfaces. On concrete, the average gait displacement per gait of both 1ML $(0.65cm,0.44cm)$ and 2ML $(0.20cm,0.15cm)$ is greater than 0ML $(0.07cm,0.06cm)$. On brick, both 1ML $(0.22cm,0.19cm)$ and 2ML $(0.13cm,0.15cm)$ outperform 0ML $(0.02cm,0.01cm)$. The relative standard deviation (standard deviation/mean) for 1ML and 2ML is lower than that for 0ML, indicating greater grip stability through improved repeatability.
@@
Examples of a single trial of each prototype row on each surface column is visualized in Fig. 12 with additional data in the accompanying video. The starting and end points are green and blue dots, and the path of traversal is a red, gradient line. On all surfaces, 1ML interacts with the environment significantly more than the baseline 0ML, resulting in greater overall movement. On every surface except for compact sand, 2ML outperforms 0ML. On concrete and compact sand, 2ML had the lowest standard deviation and on the forest floor, 1ML had the lowest standard deviation. This indicates the addition of microspine arrays also increasing the consistency and repeatability of planar locomotion with SoRos.
@@ V CONCLUSION AND FUTURE WORK
SoRos have shown immense potential with inherent conformability and adaptability to a multitude of surfaces, yet they previously lacked adequate grip stability to overcome non-uniform surfaces present outside of lab environments. Compliant microspines are one missing piece towards shrinking this real-life realization gap. We propose an elegant, compliant microspine design with a standardized soft-compliant integration technique. The stacked array configuration enables the SoRo to maintain surface interaction when extreme surface discrepancies are present. We provide results from a set of field experiments reflecting the improved performance of two microspine array configurations over a baseline SoRo on four different, ruggedized surfaces. Our results indicate that microspines are a vital technology for increasing terrain traversability in mobile SoRos. Future work includes optimizing microspine array configurations for different surfaces, performing additional field experiments, and exploring the generalizability of the design to different prototypes.

### v2

@@ IV RESULTS
The translation results of 0ML, 1ML, and 2ML are compared on the four experiment surfaces. Fig. 10 shows the average displacement and standard deviation of the three trials (60 gaits/trial = 180 gaits) for each robot configuration.
@@
The concrete surface is level with a uniform distribution of asperities. Here, both 1ML and 2ML outperform 0ML in terms of displacement. In fact, 1ML had an average displacement of 39.63cm which >30×>30$\times$> 30 × greater than travelled by 0ML, 1.32cm. 2ML traveled 14.13cm, >10×>10$\times$> 10 × as much as 0ML. 2ML had higher consistency across trials, having a standard deviation of 0.76 whereas 0ML’s standard deviation was 0.83.
@@
The next surface of partially uniform brick contains gaps in between bricks that can cause microspines to become stuck. This was observed to happen randomly across all trials. However, in every instance where this occurred, given the soft and deformable nature of the soft robot, the stuck microspine were able to wiggle free, enabling continued movement of the robots. Despite these difficulties, both 1ML (traveling 15cm) and 2ML (traveling 10.52cm) had far greater average displacement than 0ML (0.59cm), roughly 25×25$\times$25 × and 18×18$\times$18 × more, respectively. This was the only surface where 0ML had the lowest ‘gait’ standard deviation, which can be attributed to the other two robots randomly getting stuck when crossing over brick perimeters.
@@
The third surface of granular, compact sand was not entirely level and various pebbles, holes, insects, and small sticks were scattered around the testing area. This surface was difficult to overcome for all three robots as the compact sand was covered in a loose, granular top layer that was easy to become partially submerged in. Due to this, the average displacement is far lower on this surface. The two microspine limbs on 2ML seemed to dig itself deeper, resulting in average displacement of only 0.46cm and the lowest standard deviation. 1ML still outperformed 0ML with over 3 times increased displacement, traveling 2.53cm compared to 0.82cm.
@@
Figure 10: The average displacement data per prototype and surface combination. The standard deviation is indicated by error bars.
@@
Figure 11: Gait analysis. a) Every $\Delta$X and $\Delta$Y position per gait on concrete and brick. b) The average absolute $\Delta$X and $\Delta$Y displacement per gait on concrete and brick.
@@
The final surface of forest floor surface was completely unstructured and non-uniform with highly varying terrain. Specifically, the prototypes had to first overcome a large, 4” tall tree root and then move through leaves, acorns, and other tree debris. All SoRo configurations were able to successfully traverse over the cylinder-like tree root due to the conformable nature of the soft limbs, highlighted in Fig. 2. However, only the 1ML and 2ML were able to navigate through the thick tree debris after making it over the large tree root; 0ML became stuck at the base of the tree root in each of its three trials. This is exhibited in the average displacement of 13.74cm for 0ML, 23.95cm for 1ML, and 21.32cm for 2ML. The forest floor was critical for testing as it was the most unstructured of the four experimental surfaces and showcases the benefits of the soft limbs paired with the added gripping stability of the microspine array.
@@
For each of the 36 trials, the robots move for 60 gaits, resulting in a total of 2,160 gaits or 720 gaits per robot. Each gait results in relative change in position in the robot coordinate system $\Delta X,\Delta Y$. The locomotion consistency can be analyzed by examining the 180 poses on a given surface per prototype, such as those shown in Fig. 11a). The average displacement per gait ($\Delta X,\Delta Y)$ as well as the standard deviation over all gaits per prototype/surface combination is visualized in Fig. 11b). Only the uniform/partially uniform surfaces were analyzed as the gait-to-gait movement is much less consistent on the other two surfaces due to non-uniformity. On concrete, the average gait displacement per gait of both 1ML (0.65cm,0.44cm) and 2ML (0.20cm,0.15cm) is greater than 0ML (0.07cm,0.06cm). On brick, both 1ML (0.22cm,0.19cm) and 2ML (0.13cm,0.15cm) outperform 0ML (0.02cm,0.01cm). The relative standard deviation (RSD), is defined by the following equation:
$RSD=\frac{\text{standard deviation}}{\text{mean}}\times 100$
The RSD on concrete is 135.96% for 0ML, 37.05% for 1ML, and 70.55% for 2ML. On the brick surface, the RSD is 159.94% for 0ML, 54.24% for 1ML, and 89.95% for 2ML. Both 1ML and 2ML have a significantly lower RSD than 0ML, indicating the microspine array provides greater grip stability through improved repeatability.
@@
Examples of a single trial of each prototype row on each surface column is visualized in Fig. 12 with additional data in the accompanying video. The starting and end points are green and blue dots, and the path of traversal is a red, gradient line. On all surfaces, 1ML interacts with the environment significantly more than the baseline 0ML, resulting in greater overall movement. On every surface except for compact sand, 2ML outperforms 0ML.
@@ V CONCLUSION AND FUTURE WORK
SoRos have shown immense potential with inherent conformability and adaptability to a multitude of surfaces, yet they previously lacked adequate grip stability to overcome non-uniform surfaces present outside of lab environments. Improving environment interaction using compliant microspines is one of the missing piece that will facilitate shrinking of the real-life realization gap. The proposed compliant microspine design resolves the stiffness mismatch through soft-compliant integration technique. The stacked array configuration enables the SoRo to maintain surface interaction when extreme surface discrepancies are present. The results from field experiments reflect the improved performance (grip stability and repeatability) of two microspine array configurations over a baseline SoRo on four different, ruggedized surfaces. They confirm that microspines are a vital technology for increasing terrain traversability in mobile SoRos. Future work includes optimizing microspine array configurations for different surfaces, performing additional field experiments, and exploring the generalizability of the design to different prototypes.